# The Multi-Wavelength Context of Delayed Radio Emission in TDEs: Evidence for Accretion-Driven Outflows

### v1

@@ 4.2.3 Highly sub-Eddington accretion
One caveat to this picture is that while some models predict a correlation between the between the SMBH mass and the timescale required to reach the low-hard state transition (typically hundreds of days; e.g., De Colle et al. 2012; Metzger 2022), we observe no clear correlation in our data between the SMBH mass and the onset or peak timescales of the radio emission, or between the SMBH mass and the peak radio luminosity. However, as noted in Section 3.2 and C24, the exact onset time of radio emission is poorly constrained for some events. In particular, seven of our TDEs have no radio observations at $t<180$ d and others are only sparsely observed on these timescales, meaning that their prompt radio emission could have been missed entirely. If these TDEs were misclassified in our analysis, then the simple statistical tests we conduct here may not accurately reflect the underlying population. We therefore do not rule out these models immediately based on the lack of an obvious trend between SMBH mass and radio onset time, but instead choose to defer a more in-depth analysis to when a larger sample of TDEs with well-sampled radio light curves at both early and late times is available.
4.3 Origin of the prompt radio emission
@@
Finally, we briefly consider the origin of prompt radio emission in our TDEs. While various processes can delay accretion onto the SMBH relative to the $\dot{M}$ curves predicted by MOSFiT (e.g., Metzger 2022), it is more difficult to devise a mechanism by which the SMBH accretion could peak earlier than the MOSFiT prediction. This suggests that some other mechanism occurring earlier in the disruption process may be required to explain radio emission in some TDEs at early times – perhaps collision-induced outflows launched during the initial circularization of the stellar debris stream (Dai et al., 2013; Guillochon & Ramirez-Ruiz, 2015; Jiang et al., 2016; Sądowski et al., 2016; Lu & Bonnerot, 2020), or the interaction of the unbound portion of the stellar debris stream with the circumnuclear medium (Krolik et al., 2016; Yalinewich et al., 2019). However, in the latter model we would not expect any correlation with the Fundamental Plane, in contrast to our observations (Figure 5).
@@
The observed adherence of our prompt radio TDEs to the Fundamental Plane may instead indicate that a small amount of stellar material does manage to accrete onto the SMBH even at early times. However, the X-ray emission of most Prompt radio TDEs, including the canonical event ASASSN-14li (squares in Figure 5), is much softer than that of the AGN that define the Fundamental Plane, arguing against a standard XRB-like low-hard state for this early phase (Pasham & van Velzen, 2018; Guolo et al., 2023). Our other highlighted Prompt radio TDE, AT2019azh, has one of the most extensive radio and X-ray datasets of any TDE; the data included in Figure 5 span $\sim 40-800$ days post-discovery (blue-edged diamonds). Interestingly, it remains roughly consistent with the Fundamental Plane for this entire time, despite late-time changes in its X-ray hardness that suggest possible XRB-like state changes in its accretion flow (Sfaradi et al., 2022). The proximity of the Prompt radio TDEs to the Fundamental Plane could conceivably be a coincidence given the small sample size, as noted by Pasham & van Velzen (2018). Indeed, the X-ray non-detections of AT2020vwl, AT2019qiz (at early times), and ASASSN-19bt (at late times) are potentially consistent with deviations from the Fundamental Plane, but deeper observations would be required to confirm this. As the number of TDEs in the published literature with sufficiently sensitive high-cadence radio and X-ray detections at early times remains extremely small, further exploration of these possibilities will require a larger sample of promptly-observed radio TDEs and is beyond the scope of this work.
5 Conclusions
@@
  • We use the TDE modeling code MOSFiT to predict the accretion rate onto the SMBH as a function of time for each TDE and estimate the timing of likely state transitions in the accretion flow. For a few TDEs (e.g., AT2018hco, the second radio flare in AT2019dsg), the transition to a highly sub-Eddington state $\dot{M}<0.03\dot{M}_{Edd}$ is roughly coincident with the launch time inferred for the delayed radio outflow, possibly suggesting that their radio emission is related to this state transition.
  • A larger number of TDEs may have radio outflows launched coincident with the $\dot{M}\sim 0.03\dot{M}_{Edd}$ state transition if this transition is delayed by several hundred days relative to the MOSFiT prediction. This is also supported by the X-ray behavior of these TDEs, which are consistent with the Fundamental Plane of black hole activity at late times. For ASASSN-14ae, this transition may happen as late as $\sim 2000$ days post-discovery.
  • Other TDEs (e.g., ASASSN-15oi, AT2018hyz) are much more radio-bright/X-ray dim than is typical for black holes accreting at a few percent of the Eddington rate, as seen by their increasing distance off the BH fundamental plane with time (Figure 5). We suggest that this sub-group of Delayed radio TDEs is powered instead by jets or outflows launched during a period of super-Eddington accretion. For these TDEs, the lack of bright radio emission at early times then implies either that the radio emission is powered by a promptly-launched off-axis jet initially suppressed by relativistic beaming, or that super-Eddington accretion is delayed by several hundred days relative to peak optical light.
  • TDEs with delayed radio emission may be identifiable at early times by their larger optical photospheric radii and lack of He emission at peak. If this indicates slower debris circularization and delayed-onset accretion as previously suggested by (Nicholl et al., 2022), the super-Eddington accretion phase would also be later than MOSFiT’s prediction and some delayed radio TDEs could be powered by non-relativistic outflows launched during this phase. If this correlation between TDE optical and radio properties is confirmed in a larger sample of TDEs, it may provide an important tool to predict which TDEs are likely to brighten in the radio at late times, enabling more efficient use of limited follow up resources.
@@
As larger samples of TDEs are discovered and the time cost of intensive targeted follow up becomes prohibitive, all-sky radio surveys will play an increasingly important role in systematically monitoring the full TDE population. Our independent VLASS discovery of the dramatic rebrightening of ASASSN-15oi (Hajela et al., 2024) demonstrates the value of unbiased radio surveys for uncovering unexpected behavior – it is unlikely that late-rising radio emission in TDEs would have been explored at all, without the initial VLASS detection of this TDE. Our few VLASS detections (Appendix A) confirm that $2-6$ years post-disruption may be a particularly fruitful time to search for late-rising low-frequency radio emission in optically selected TDEs.
@@
All-sky radio surveys are now also uncovering new nuclear transients that may be TDEs missed at other wavelengths (e.g., Anderson et al. 2020; Ravi et al. 2022; Somalwar et al. 2023b, a; Dykaar et al. 2024). Our results suggest that deep late-time X-ray observations of these ambiguous radio transients may be one way to gain insight even when optical and X-ray observations extending to several years prior to the radio discovery are not available. While VLASS is the deepest radio survey of the northern sky conducted to date, many of the TDE outflows observed in the targeted radio observations of our sample are too faint to be detected in VLASS for all but the nearest TDEs, suggesting that VLASS’s limited sensitivity, frequency coverage, and time sampling may still be inadequate to fully characterize this previously unappreciated late phase of TDEs’ radio evolution. The planned ngVLA combined with wide-field radio surveys with the SKA will greatly increase our sensitivity to both low-luminosity radio outflows from TDEs and to distant jetted TDEs (Donnarumma et al., 2015; van Velzen et al., 2018), providing critical support for the large numbers of higher-redshift TDEs expected from Rubin Observatory and other upcoming optical and multi-wavelength surveys.
@@
We thank Sjoert van Velzen, Adelle Goodwin, James Miller-Jones, and Nayana AJ for useful conversations. The National Radio Astronomy Observatory (NRAO) is a facility of the U.S. National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This research has made use of the CIRADA cutout service at URL cutouts.cirada.ca, operated by the Canadian Initiative for Radio Astronomy Data Analysis (CIRADA). CIRADA is funded by a grant from the Canada Foundation for Innovation 2017 Innovation Fund (Project 35999), as well as by the Provinces of Ontario, British Columbia, Alberta, Manitoba and Quebec, in collaboration with the National Research Council of Canada, the US National Radio Astronomy Observatory and Australia’s Commonwealth Scientific and Industrial Research Organisation. In addition to Swift observations proposed by us, we have also used the Swift archive at https://swift.gsfc.nasa.gov/archive/. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. KDA and CTC acknowledge support provided by the NSF through award SOSPA9-007 from the NRAO and award AST-2307668. KDA gratefully acknowledges support from the Alfred P. Sloan Foundation. NF acknowledges support from the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2137419. BDM was supported in part by the National Science Foundation (grant No. AST-2009255). The Flatiron Institute is supported by the Simons Foundation. MN is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 948381) and by UK Space Agency Grant No. ST/Y000692/1. ER-R was supported in part by the Heising-Simons Foundation, NSF (AST-2150255 and AST-2307710), Swift (80NSSC21K1409,80NSSC19K1391) and Chandra (22-0142). FDC acknowledges support from the UNAM-PAPIIT grant IN113424. This research was supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). RC and RM acknowledge the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452, for its hospitality while this work was completed. , CASA (McMullin et al., 2007), pwkit (Williams et al., 2017)
@@ Appendix A VLASS search for late-rising radio emission from known TDEs
The advent of the VLASS, the most sensitive radio survey of the full northern sky to date, provides an exciting new opportunity to search for late-rising radio emission from known TDE candidates. VLASS began on 2017 September 7 and is using the VLA to map 80% of the radio sky in three epochs spanning seven years (Lacy et al., 2020). All VLASS observations are taken at $2-4$ GHz (S-band) and the typical rms sensitivity of the single-epoch images is $\sim 100$ $\mu$Jy.
Given the newly-recognized prevalence of long-lasting radio emission in TDEs, all-sky radio surveys also open up a complementary method of discovering new TDEs in the nearby Universe. Blind radio searches with VLASS and radio surveys with the SKA precursors are now finding dozens of nuclear transients that may be TDEs, but interpretation is difficult without multi-frequency radio follow up (e.g., Anderson et al. 2020; Ravi et al. 2022; Somalwar et al. 2023b, a; Dykaar et al. 2024). The rates and higher average radio luminosities of transients discovered in these blind searches suggest that not all of them are drawn from the same underlying population as the TDEs studied in this work, but it is not clear if this reflects a true dichotomy or a bias in host environment – for example, TDEs in dense gas and dust-rich environments may produce stronger radio emission (Generozov et al., 2017), but may have weaker observed optical and X-ray emission due to extinction. VLASS allows us to easily compare the radio properties of optical TDEs to those of other nuclear transients.
@@
We cross-matched the locations of these 91 TDE candidates with our Source Extractor catalog of possible VLASS source detections, publicly available here: 10.5281/zenodo.4895112. The creation of this catalog is described in Stroh et al. (2021). For this search, we utilized the Epoch 1 quick-look images provided by NRAO. [^19] For additional information about the VLASS quick-look images and their limitations, see https://science.nrao.edu/vlass/data-access/vlass-epoch-1-quick-look-users-guide. The first half of the data (VLASS1.1) was collected between September 2017 and February 2018 and suffers from additional uncertainties in the absolute flux calibration scale; the peak flux densities of sources fainter than 1 Jy are systematically low by 15%, with an additional systematic scatter of ±$\sim\pm 8$%. The second half of the data (VLASS1.2) was collected between February and July 2019; the total flux densities of faint sources in these data are low by $\sim 8$% with a scatter of $\sim 5\%$. As our primary goal is simply to determine the possible presence of transient radio emission, these calibration uncertainties are acceptable for our purposes.
@@
We identified 5 TDEs located within 5 arcsec of a VLASS source detected with a signal-to-noise ratio $>5$ in our catalog, which are presented in Table 3. We use the imfit command within CASA to fit an elliptical Gaussian to each TDE and take the peak brightness to be equal to the flux density, as expected for an unresolved point source. Interestingly, each of these TDEs was discovered via a different method: NGC 5905 was one of the first TDE candidates ever reported and was identified as an X-ray flare from the center of a non-AGN galaxy (Bade et al., 1996); Swift J1644+57 was discovered as a long-lasting, luminous gamma-ray and X-ray transient (Bloom et al., 2011; Burrows et al., 2011; Levan et al., 2011); ASASSN-15oi was discovered in an optical survey and classified as a TDE based on its nuclear location, photometric evolution, and spectroscopic properties (Holoien et al., 2016a); J100933 was discovered as a mid-infrared flare (Wang et al., 2018); and F01004-2237 was identified based on changes to its optical spectrum observed serendipitously during a survey of nearby ultra-luminous infrared galaxies (Tadhunter et al., 2017). Two of these five objects, Swift J1644+57 and ASASSN-15oi, have well-studied radio counterparts previously reported in the literature (Zauderer et al., 2011; Berger et al., 2012; Wiersema et al., 2012; Zauderer et al., 2013; Yang et al., 2016; Eftekhari et al., 2018; Cendes et al., 2021b; Horesh et al., 2021a; Hajela et al., 2024). Swift J1644+57’s radio emission, first detected within days of its discovery, is well-explained as synchrotron emission from an initially relativistic jet viewed on-axis (or slightly off-axis; see Beniamini et al. 2023). Its detection in VLASS epochs 1 and 2 was previously reported by Cendes et al. (2021b). The epoch 1 VLASS detection of ASASSN-15oi, also reported by Horesh et al. (2021a) and Hajela et al. (2024), was the first evidence for a dramatic second brightening of its radio emission, providing the initial impetus for the deeper late-time observations of other TDEs presented in C24 and this work. NGC 5905 was the first TDE with a follow-up radio observation to search for jet emission. A VLA observation $\sim$ 6 years after the X-ray peak provided an upper limit of $<0.15$ mJy at 8.46 GHz for a central point source (Komossa & Dahlem, 2001). It was also observed as part of the NRAO VLA Sky Survey (NVSS). The NVSS detection of extended radio emission of NGC 5905 (21 mJy at 1.4 GHz) was found to be consistent with the starburst properties of the host galaxy and consistent with pre-outburst extended radio emission (Komossa, 2002; Raichur et al., 2015). The VLASS is less sensitive to extended emission due to the survey’s higher resolution, but the VLASS flux densities of NGC 5905 are approximately constant in time and are also likely consistent with unrelated emission from the host galaxy. We include NGC 5905’s VLASS flux density in Table 3 for completeness but do not investigate this source further. We also choose not to further observe F01004-2237, as the nature of this source is disputed. It was initially proposed to be a TDE (Tadhunter et al., 2017), later re-classified as an unusual AGN Bowen fluorescence flare (Trakhtenbrot et al., 2019), and finally recently suggested to be a repeating partial TDE (or two independent TDEs) after a second optical flare was observed in September 2021 (Sun et al., 2024). The continued activity of this source motivates future multi-wavelength monitoring to distinguish between the AGN and repeated or partial TDE scenarios.
@@
We triggered additional multi-frequency follow up observations of the two most promising radio TDE candidates in our sample, ASASSN-15oi and J100933. We observed these targets with the VLA on 2020 May 20 (program 20A-492, PI: Alexander) and 2021 June 10 (program 21A-303, PI: Hajela). Observations were conducted at $1-12$ GHz (L, S, C, and X bands) when the VLA was in its C configuration. The data were reduced in CASA using standard procedures and flux densities were measured using the imtool fitsrc command within the pwkit package (Williams et al., 2017). We have verified that this gives similar results to measuring the flux densities with imfit. The follow up observations of ASASSN-15oi are presented in Hajela et al. (2024). We include the measured flux densities of J100933 in Table 3 together with their statistical errors. The systematic uncertainty on the flux calibration for these observations is 5% at all frequencies. In our 2021 observations, J100933 exhibits a rising spectrum with a peak around 9 GHz. We find that the flux density of J100933 at 3 GHz increased slightly between the time of the VLASS epoch 1 detection and our follow up observations two years later and then remained approximately constant for the next three years. Given the additional flux calibration uncertainties in the quick look images for epoch 1 of VLASS, this is roughly consistent with no significant temporal evolution over the full 5-year period. We thus conclude that the radio emission in this target is likely unrelated to the transient, although longer-term multi-frequency monitoring may be needed to confirm this due to the age of the TDE ($>11$ yr).
@@
We also visually inspected the coordinates of each of the 91 TDEs in our sample in VLASS epochs 2.1, 2.2, and 3.1 quicklook data using the CIRADA cutout server to see if any of them appeared in the radio after VLASS epoch 1. We find only one new TDE undetected in VLASS epoch 1 that “turns on” in subsequent data: AT2018hyz, which is not detected in epochs 1 or 2 but appears as a bright point source in VLASS epoch 3. The VLASS data for AT2018hyz have already been presented in Cendes et al. (2022b) and C24; we therefore do not repeat them here. We also continue to detect the five sources visible in our Epoch 1 search. NGC 5905, F01004-2237, and J100933 exhibit roughly constant radio brightness across all VLASS epochs, while Sw J1644+57 and ASASSN-15oi slowly fade. We include the flux densities of these five TDEs in all available VLASS epochs in Table 3.

### v2

@@ IV.2.3 Highly sub-Eddington accretion
One caveat to this picture is that while some models predict a correlation between the between the SMBH mass and the timescale required to reach the low-hard state transition (typically hundreds of days; e.g., De Colle et al. 2012; Metzger 2022), we observe no clear correlation in our data between the SMBH mass and the onset or peak timescales of the radio emission, or between the SMBH mass and the peak radio luminosity. However, as noted in Section III.2 and C24, the exact onset time of radio emission is poorly constrained for some events. In particular, seven of our TDEs have no radio observations at $t<180$ d and others are only sparsely observed on these timescales, meaning that their prompt radio emission could have been missed entirely. If these TDEs were misclassified in our analysis, then the simple statistical tests we conduct here may not accurately reflect the underlying population. We therefore do not rule out these models immediately based on the lack of an obvious trend between SMBH mass and radio onset time, but instead choose to defer a more in-depth analysis to when a larger sample of TDEs with well-sampled radio light curves at both early and late times is available.
IV.3 Origin of the prompt radio emission
@@
Finally, we briefly consider the origin of prompt radio emission in our TDEs. While various processes can delay accretion onto the SMBH relative to the $\dot{M}$ curves predicted by MOSFiT (e.g., Metzger 2022), it is more difficult to devise a mechanism by which the SMBH accretion could peak earlier than the MOSFiT prediction. This suggests that some other mechanism occurring earlier in the disruption process may be required to explain radio emission in some TDEs at early times – perhaps collision-induced outflows launched during the initial circularization of the stellar debris stream (Dai et al., 2013; Guillochon and Ramirez-Ruiz, 2015; Jiang et al., 2016; Sądowski et al., 2016; Lu and Bonnerot, 2020), or the interaction of the unbound portion of the stellar debris stream with the circumnuclear medium (Krolik et al., 2016; Yalinewich et al., 2019). Both of these models can explain the range of radio luminosities observed in TDEs at early times. However, in the latter model we would not expect any correlation with the Fundamental Plane, in contrast to our observations (Figure 5).
@@
The observed adherence of our prompt radio TDEs to the Fundamental Plane may instead indicate that a small amount of stellar material does manage to accrete onto the SMBH even at early times. However, the X-ray emission of most Prompt radio TDEs, including the canonical event ASASSN-14li (squares in Figure 5), is much softer than that of the AGN that define the Fundamental Plane, arguing against a standard XRB-like low-hard state for this early phase (Pasham and van Velzen, 2018; Guolo et al., 2024). Our other highlighted Prompt radio TDE, AT2019azh, has one of the most extensive radio and X-ray datasets of any TDE; the data included in Figure 5 span $\sim 40-800$ days post-discovery (blue-edged diamonds). Interestingly, it remains roughly consistent with the Fundamental Plane for this entire time, despite late-time changes in its X-ray hardness that suggest possible XRB-like state changes in its accretion flow (Sfaradi et al., 2022). The proximity of the Prompt radio TDEs to the Fundamental Plane could conceivably be a coincidence given the small sample size, as noted by Pasham and van Velzen (2018). Indeed, the X-ray non-detections of AT2020vwl, AT2019qiz (at early times), and ASASSN-19bt (at late times) are potentially consistent with deviations from the Fundamental Plane, but deeper observations would be required to confirm this. As the number of TDEs in the published literature with sufficiently sensitive high-cadence radio and X-ray detections at early times remains extremely small, further exploration of these possibilities will require a larger sample of promptly-observed radio TDEs and is beyond the scope of this work.
V Conclusions
@@
  • We use the TDE modeling code MOSFiT to predict the accretion rate onto the SMBH as a function of time for each TDE and estimate the timing of likely state transitions in the accretion flow. For at least 3, and up to 10 TDEs (e.g., AT2018hco, the second radio flare in AT2019dsg), the transition to a highly sub-Eddington state $\dot{M}<0.03\,\dot{M}_{\rm Edd}$ is roughly coincident with the launch time inferred for the delayed radio outflow, possibly suggesting that their radio emission is related to this state transition.
  • A further 3-6 TDEs may have radio outflows launched coincident with the $\dot{M}\sim 0.03\,\dot{M}_{\rm Edd}$ state transition if this transition is delayed by several hundred days relative to the MOSFiT prediction. This is also supported by the X-ray behavior of these TDEs, which are consistent with the Fundamental Plane of black hole activity at late times. For ASASSN-14ae, this transition may happen as late as $\sim 2000$ days post-discovery.
  • Other TDEs (ASASSN-15oi, AT2018hyz) are much more radio-bright/X-ray dim than is typical for black holes accreting at a few percent of the Eddington rate, as seen by their increasing distance off the BH fundamental plane with time (Figure 5). We suggest that this sub-group of delayed radio TDEs is powered instead by jets or outflows launched during a period of super-Eddington accretion. For these TDEs, the lack of bright radio emission at early times then implies either that the radio emission is powered by a promptly-launched off-axis jet initially suppressed by relativistic beaming, or that super-Eddington accretion is delayed by several hundred days relative to peak optical light.
  • TDEs with delayed radio emission may be identifiable at early times by their larger optical photospheric radii and lack of He emission at peak. If this indicates slower debris circularization and delayed-onset accretion as previously suggested by (Nicholl et al., 2022), the super-Eddington accretion phase would also be later than MOSFiT’s prediction and some delayed radio TDEs could be powered by non-relativistic outflows launched during this phase. A super-Eddington phase delayed by up to several hundred days relative to MOSFiT’s prediction could explain the radio emission seen in 11/12 of our delayed radio TDEs plus the second flare in AT2019dsg; the exception is ASASSN-14ae. If this correlation between TDE optical and radio properties is confirmed in a larger sample of TDEs, it may provide an important tool to predict which TDEs are likely to brighten in the radio at late times, enabling more efficient use of limited follow up resources.
@@
As larger samples of TDEs are discovered and the time cost of intensive targeted follow up becomes prohibitive, all-sky radio surveys will play an increasingly important role in systematically monitoring the full TDE population. Our independent VLASS discovery of the dramatic rebrightening of ASASSN-15oi (Hajela et al., 2025) demonstrates the value of unbiased radio surveys for uncovering unexpected behavior — it is unlikely that late-rising radio emission in TDEs would have been explored at all, without the initial VLASS detection of this TDE. Our few VLASS detections (Appendix A) confirm that 262-6 years post-disruption may be a particularly fruitful time to search for late-rising low-frequency radio emission in optically selected TDEs.
@@
All-sky radio surveys are now also uncovering new nuclear transients that may be TDEs missed at other wavelengths (e.g., Anderson et al. 2020; Ravi et al. 2022; Somalwar et al. 2023, 2025; Dykaar et al. 2024). Our results suggest that deep late-time X-ray observations of these ambiguous radio transients may be one way to gain insight even when optical and X-ray observations extending to several years prior to the radio discovery are not available. While VLASS is the deepest radio survey of the northern sky conducted to date, many of the TDE outflows observed in the targeted radio observations of our sample are too faint to be detected in VLASS for all but the nearest TDEs, suggesting that VLASS’s limited sensitivity, frequency coverage, and time sampling may still be inadequate to fully characterize this previously unappreciated late phase of TDEs’ radio evolution. The planned ngVLA combined with wide-field radio surveys with the SKA will greatly increase our sensitivity to both low-luminosity radio outflows from TDEs and to distant jetted TDEs (Donnarumma et al., 2015; van Velzen et al., 2018), providing critical support for the large numbers of higher-redshift TDEs expected from Rubin Observatory and other upcoming optical and multi-wavelength surveys.
@@
We thank Sjoert van Velzen, Adelle Goodwin, James Miller-Jones, and Nayana AJ for useful conversations. We also thank the anonymous referee for helpful comments that have improved this work. The National Radio Astronomy Observatory (NRAO) is a facility of the U.S. National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This research has made use of the CIRADA cutout service at URL cutouts.cirada.ca, operated by the Canadian Initiative for Radio Astronomy Data Analysis (CIRADA). CIRADA is funded by a grant from the Canada Foundation for Innovation 2017 Innovation Fund (Project 35999), as well as by the Provinces of Ontario, British Columbia, Alberta, Manitoba and Quebec, in collaboration with the National Research Council of Canada, the US National Radio Astronomy Observatory and Australia’s Commonwealth Scientific and Industrial Research Organisation. In addition to Swift observations proposed by us, we have also used the Swift archive at https://swift.gsfc.nasa.gov/archive/. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. KDA and CTC acknowledge support provided by the NSF through award SOSPA9-007 from the NRAO and award AST-2307668. KDA gratefully acknowledges support from the Alfred P. Sloan Foundation. NF acknowledges support from the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2137419. BDM was supported in part by the National Science Foundation (grant No. AST-2009255). The Flatiron Institute is supported by the Simons Foundation. MN is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 948381) and by UK Space Agency Grant No. ST/Y000692/1. ER-R was supported in part by the Heising-Simons Foundation, NSF (AST-2150255 and AST-2307710), Swift (80NSSC21K1409,80NSSC19K1391) and Chandra (22-0142). FDC acknowledges support from the UNAM-PAPIIT grant IN113424. This research was supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). RC and RM acknowledge the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452, for its hospitality while this work was completed.
, CASA (McMullin et al., 2007), pwkit (Williams et al., 2017)
@@ Appendix A VLASS search for late-rising radio emission from known TDEs
[…]
Given the newly-recognized prevalence of long-lasting radio emission in TDEs, all-sky radio surveys also open up a complementary method of discovering new TDEs in the nearby Universe. Blind radio searches with VLASS and radio surveys with the SKA precursors are now finding dozens of nuclear transients that may be TDEs, but interpretation is difficult without multi-frequency radio follow up (e.g., Anderson et al. 2020; Ravi et al. 2022; Somalwar et al. 2023, 2025; Dykaar et al. 2024). The rates and higher average radio luminosities of transients discovered in these blind searches suggest that not all of them are drawn from the same underlying population as the TDEs studied in this work, but it is not clear if this reflects a true dichotomy or a bias in host environment – for example, TDEs in dense gas and dust-rich environments may produce stronger radio emission (Generozov et al., 2017), but may have weaker observed optical and X-ray emission due to extinction. VLASS allows us to easily compare the radio properties of optical TDEs to those of other nuclear transients.
@@
We cross-matched the locations of these 91 TDE candidates with our Source Extractor catalog of possible VLASS source detections, publicly available on Zenodo (Stroh et al., 2021a). The creation of this catalog is described in Stroh et al. (2021b). For this search, we utilized the Epoch 1 quick-look images provided by NRAO. [^21] For additional information about the VLASS quick-look images and their limitations, see https://science.nrao.edu/vlass/data-access/vlass-epoch-1-quick-look-users-guide. The first half of the data (VLASS1.1) was collected between September 2017 and February 2018 and suffers from additional uncertainties in the absolute flux calibration scale; the peak flux densities of sources fainter than 1 Jy are systematically low by 15%, with an additional systematic scatter of $\sim\pm 8$%. The second half of the data (VLASS1.2) was collected between February and July 2019; the total flux densities of faint sources in these data are low by $\sim 8$% with a scatter of $\sim 5\%$. As our primary goal is simply to determine the possible presence of transient radio emission, these calibration uncertainties are acceptable for our purposes.
@@
We identified 5 TDEs located within 5 arcsec of a VLASS source detected with a signal-to-noise ratio >5>5 in our catalog, which are presented in Table 5. We use the imfit command within CASA to fit an elliptical Gaussian to each TDE and take the peak brightness to be equal to the flux density, as expected for an unresolved point source. Interestingly, each of these TDEs was discovered via a different method: NGC 5905 was one of the first TDE candidates ever reported and was identified as an X-ray flare from the center of a non-AGN galaxy (Bade et al., 1996); Swift J1644+57 was discovered as a long-lasting, luminous gamma-ray and X-ray transient (Bloom et al., 2011; Burrows et al., 2011; Levan et al., 2011); ASASSN-15oi was discovered in an optical survey and classified as a TDE based on its nuclear location, photometric evolution, and spectroscopic properties (Holoien et al., 2016a); J100933 was discovered as a mid-infrared flare (Wang et al., 2018); and F01004-2237 was identified based on changes to its optical spectrum observed serendipitously during a survey of nearby ultra-luminous infrared galaxies (Tadhunter et al., 2017). Two of these five objects, Swift J1644+57 and ASASSN-15oi, have well-studied radio counterparts previously reported in the literature (Zauderer et al., 2011; Berger et al., 2012; Wiersema et al., 2012; Zauderer et al., 2013; Yang et al., 2016; Eftekhari et al., 2018; Cendes et al., 2021b; Horesh et al., 2021a; Hajela et al., 2025). Swift J1644+57’s radio emission, first detected within days of its discovery, is well-explained as synchrotron emission from an initially relativistic jet viewed on-axis (or slightly off-axis; see Beniamini et al. 2023). Its detection in VLASS epochs 1 and 2 was previously reported by Cendes et al. (2021b). The epoch 1 VLASS detection of ASASSN-15oi, also reported by Horesh et al. (2021a) and Hajela et al. (2025), was the first evidence for a dramatic second brightening of its radio emission, providing the initial impetus for the deeper late-time observations of other TDEs presented in C24 and this work. NGC 5905 was the first TDE with a follow-up radio observation to search for jet emission. A VLA observation $\sim$ 6 years after the X-ray peak provided an upper limit of <0.15<0.15 mJy at 8.46 GHz for a central point source (Komossa and Dahlem, 2001). It was also observed as part of the NRAO VLA Sky Survey (NVSS). The NVSS detection of extended radio emission of NGC 5905 (21 mJy at 1.4 GHz) was found to be consistent with the starburst properties of the host galaxy and consistent with pre-outburst extended radio emission (Komossa, 2002; Raichur et al., 2015). The VLASS is less sensitive to extended emission due to the survey’s higher resolution, but the VLASS flux densities of NGC 5905 are approximately constant in time and are also likely consistent with unrelated emission from the host galaxy. We include NGC 5905’s VLASS flux density in Table 5 for completeness but do not investigate this source further. We also choose not to further observe F01004-2237, as the nature of this source is disputed. It was initially proposed to be a TDE (Tadhunter et al., 2017), later re-classified as an unusual AGN Bowen fluorescence flare (Trakhtenbrot et al., 2019), and finally recently suggested to be a repeating partial TDE (or two independent TDEs) after a second optical flare was observed in September 2021 (Sun et al., 2024). The continued activity of this source motivates future multi-wavelength monitoring to distinguish between the AGN and repeated or partial TDE scenarios.
@@
We triggered additional multi-frequency follow up observations of the two most promising radio TDE candidates in our sample, ASASSN-15oi and J100933. We observed these targets with the VLA on 2020 May 20 (program 20A-492, PI: Alexander) and 2021 June 10 (program 21A-303, PI: Hajela). Observations were conducted at 1121-12 GHz (L, S, C, and X bands) when the VLA was in its C configuration. The data were reduced in CASA using standard procedures and flux densities were measured using the imtool fitsrc command within the pwkit package (Williams et al., 2017). We have verified that this gives similar results to measuring the flux densities with imfit. The follow up observations of ASASSN-15oi are presented in Hajela et al. (2025). We include the measured flux densities of J100933 in Table 5 together with their statistical errors. The systematic uncertainty on the flux calibration for these observations is 5% at all frequencies. In our 2021 observations, J100933 exhibits a rising spectrum with a peak around 9 GHz. We find that the flux density of J100933 at 3 GHz increased slightly between the time of the VLASS epoch 1 detection and our follow up observations two years later and then remained approximately constant for the next three years. Given the additional flux calibration uncertainties in the quick look images for epoch 1 of VLASS, this is roughly consistent with no significant temporal evolution over the full 5-year period. We thus conclude that the radio emission in this target is likely unrelated to the transient, although longer-term multi-frequency monitoring may be needed to confirm this due to the age of the TDE (>11>11 yr).
@@
We also visually inspected the coordinates of each of the 91 TDEs in our sample in VLASS epochs 2.1, 2.2, and 3.1 quicklook data using the CIRADA cutout server to see if any of them appeared in the radio after VLASS epoch 1. We find only one new TDE undetected in VLASS epoch 1 that “turns on” in subsequent data: AT2018hyz, which is not detected in epochs 1 or 2 but appears as a bright point source in VLASS epoch 3. The VLASS data for AT2018hyz have already been presented in Cendes et al. (2022a) and C24; we therefore do not repeat them here. We also continue to detect the five sources visible in our Epoch 1 search. NGC 5905, F01004-2237, and J100933 exhibit roughly constant radio brightness across all VLASS epochs, while Sw J1644+57 and ASASSN-15oi slowly fade. We include the flux densities of these five TDEs in all available VLASS epochs in Table 5.